The lease of Theme Park, Inc., is about to expire. Management must decide whether to renew the lease for another 10 years or to relocate near the site of a proposed motel. The town planning board is currently debating the merits of granting approval to the motel. A consultant has estimated the net present value of Theme Park’s two alternatives under each state of nature as shown below. Suppose that the management of Theme Park, Inc., has decided that there is a 0.23 probability that the motel’s application will be approved.

Options Motel
Approved
Motel
Rejected
Renew $ 430,000 $ 4,075,000
Relocate 2,075,000 130,000


a-1.
If management uses maximum expected monetary value as the decision criterion, calculate expected monetary value for the alternatives "Renew" and "Relocate"? (Omit the "$" sign in your response.)

Alternative Expected Value
Renew $
Relocate $


a-2. Which alternative should it choose?

  • Relocate

  • Renew lease



c. If management has been offered the option of a temporary lease while the town planning board considers the motel’s application, would you advise management to sign the lease? The lease will cost $23,000. (Omit the "$" sign in your response.)

  (Click to select)   No   Yes  because the cost is   (Click to select)   greater   less  than EVPI of $  .
Answer #1

a1

NPV of renew = 430000x0.23+4075000x0.77 = 98900+3137750 =3236650

NPV of relocate =2075000x0.23+130000x0.77 =477250+100100 =577350

a2 Renewal is the more favourable option as the NPV is higher in this case

c

EVPI = EV with PI - EV without PI

= (0.23x2075000+0.77x4075000) -3236650

= (477250+3177750) -3236650

= 418350

The deal should be signed as the cost of 23000 is less than EVPI of 418350